Use the formula A=P(1+r/n)^nt to solve the compound interest problem
Find how long it takes for $1700 to double if it is invested at 6% interested compounded monthly

Answer by richwmiller(17219)

A=P(1+r/n)^nt
In the formula,
A is the amount of money in the savings account, 3400
P is the principle, 1700
r is the interest rate, 6% or .06
t is the time that the money is in the account, ---We are trying to find this.
n is the number of times the money is compounded per year. 12
A=P(1+r/n)^nt
3400=1700(1+.06/12)^12t